OpenStudy (cindy):

simplify 2(1-d) - (2d+1) (2-2d) (2d+1) 2-4d+1 = 3-4d is that correct

OpenStudy (anonymous):

\[(2-2d)-(2d+1)=2-1-4d=1-4d\]

OpenStudy (cindy):

so th ne i wrote dwn s wrong

OpenStudy (anonymous):

Yes. You have made a mistake in moving from your second step to your third. You've got the coefficient of d correct but you need to realise that subtracting (2d+1) is the same as subtracting 2d and then subtracting 1.

OpenStudy (cindy):

can yu show me the steps

OpenStudy (anonymous):

\[2(1-d)-(2d+1)=(2-2d)-(2d+1)=(2-2d)-2d-1=1-4d\]